# Periodic Asynchrony: An Effective Method for Accelerating Reinforcement Learning

### v2

@@ 5 Implementation
The implementation in this study is based on the PyTorch training framework, with the training component using a 3D-parallel distributed architecture. Parts of the code from Megatron-Core v0.8.0 are incorporated.On the NPU platform, additional components from MindSpeed v2.1.0_core_r0.8.0, torch_npu, and several optimized operators included in torch_npu—such as npu_fusion_attention, an accelerated attention kernel supporting custom masks, which is designed as a counterpart to flash attention [dao2023flashattention]—are also employed.
For the inference component, vLLM is used as the engine, and on the NPU platform, ascend_vLLM 0.9.2.rc1 is adopted.
As in fine-tuning, all samples in the reinforcement learning training process are computed with their native sequence lengths for both forward and backward passes, without any additional padding; that is, the system uses dynamic-length training.
@@ 6 Experiments
This study primarily conducts comparative experiments on mathematical reasoning tasks. The reinforcement learning frameworks included in the comparison are: MindSpeed-RL, which adopts a training–inference shared-resource architecture; the synchronous reinforcement learning version implemented in this work under a training–inference decoupled design; and the asynchronous reinforcement learning version proposed in this study.
The main objective is to experimentally validate the performance between periodic asynchronous and synchronous training, as well as to compare the proposed periodic asynchronous system with other mainstream synchronous reinforcement learning frameworks on the same platform, while also evaluating the scalability of our periodic asynchronous system. The primary evaluation metric is the end-to-end training throughput, such as the average number of tokens trained per second per device (TPSPD). This is calculated as the total number of tokens processed over several steps divided by the total elapsed time and then divided by the number of devices used. Since no modifications were made to the reinforcement learning algorithm, accuracy metrics are reported only for reference, primarily to verify the correctness of algorithm reproduction.
The experiments in this study were conducted on air-cooled Ascend-910B NPUs, with each node equipped with eight NPUs, each providing 64 GB of memory. The intra-node interconnect bandwidth is 196 GB/s, while the inter-node bandwidth is 100 Gb/s.
@@ 6.1 Model and Environment Configuration
The models compared in this section are Qwen3-8B  [qwen3technicalreport] and DeepSeek-R1-Distill-Qwen-32B  [deepseekai2025deepseekr1incentivizingreasoningcapability] (since the current version of Mindspeed-RL does not yet support reinforcement learning training for Qwen3-32B, we instead use DeepSeek-R1-Distill-Qwen-32B as the training model.). The training data comes from the publicly available math problem dataset DeepScaler [deepscaler2025]. All reinforcement learning experiments use the GRPO algorithm. The accuracy test set is the AIME24 [aime24].
@@ 6.2 Comparison between Asynchronous and Synchronous Training
In the first set of experiments, Qwen3-8B was used as the training model, with a context length of 16k. A total of 16 NPUs were used. For the synchronous and asynchronous training configurations implemented in this work under the training–inference decoupled architecture, 8 NPUs were allocated to launch a single training instance with tensor parallelism ($\text{TP}=8$) and pipeline parallelism ($\text{PP}=1$). The remaining 8 NPUs were used to launch four inference instances with $\text{TP}=2$. For MindSpeed-RL, the training configuration used 16 NPUs to launch two training instances ($\text{TP}=8$, $\text{PP}=1$), while all 16 NPUs were also used to launch eight inference instances ($\text{TP}=2$).
Table 1: Accuracy and training efficiency comparison results on Qwen3-8B. Each framework was trained for 100 steps, and the total time consumed and the total number of tokens trained were extracted from the logs.
Framework GRPO Steps AIME24 Trained Tokens Training Hours NPUs Tokens/Sec/Device
Baseline 0 0.725 0 0 0 -
MindSpeed-RL 100 0.733 706.355M 198.942 16 61.641
Sync (ours) 100 0.746 827.306M 143.678 16 99.966
Async (ours) 100 0.758 917.609M 82.861 16 192.259
@@
In terms of TPSPD, the training speed ratio among the three implementations was 1:1.62:3.12. For accuracy evaluation, we adopted a rule-based method, where the final answer predicted by the policy model is considered correct if it can be correctly extracted and matches the ground-truth answer in the dataset; otherwise, it is considered incorrect. The tests were conducted using the inference engine vLLM 0.9.1.rc, with an output length limit of 32k. Each test sample was evaluated 8 times, with inference parameters set to temperature = 0.7, top_p = 0.95, and top_k = 20. In total, 240 test problems were evaluated. It is evident that the models trained by the three frameworks achieved similar accuracy on the AIME24 benchmark, with variations within normal random fluctuations, supporting the equivalence in effectiveness between the asynchronous and synchronous methods implemented in this work.
@@
Figure 5: Training step-wise average reward score comparison
We also monitored the reward scores throughout the training process of the three frameworks, as depicted in Figure 5. The reward score trajectories of our synchronous and asynchronous methods are nearly indistinguishable, further reinforcing the assertion that the proposed asynchronous approach is computationally equivalent to its synchronous counterpart. Both curves show a slight upward trend, which can likely be attributed to the baseline model having already been pre-trained on similar data. Notably, in the initial phases of training, the reward mean curve for Mindspeed-RL diverges significantly from that of our approach. This disparity may arise from differences in system architecture. Additionally, it is important to note that, in contrast to our environment, Mindspeed-RL utilizes vLLM version 0.9.1 and vllm_ascend version 0.9.0rc2.
The model used in the second set of experiments is DeepSeek-R1-Distill-Qwen-32B, with a context length of 16k, and only two experiments are conducted. For our asynchronous framework, the configuration is as follows: 32 NPUs are used to launch a single training instance with $\text{TP}=8$ and $\text{PP}=4$, while the remaining 16 NPUs are used to launch four inference instances with $\text{TP}=4$. In contrast, for the Mindspeed-RL configuration, 64 NPUs are used to launch two training instances with $\text{TP}=8$ and $\text{PP}=4$, and simultaneously, the same 64 NPUs are used to launch 16 inference instances with $\text{TP}=4$.
Table 2: Accuracy and training efficiency comparison results on DeepSeek-R1-Distill-Qwen-32B. Each framework was trained for 20 steps, and the total time consumed and the total number of tokens trained were extracted from the logs.
Framework GRPO Steps AIME24 Trained Tokens Training Hours NPUs Tokens/Sec/Device
Baseline 0 0.726 0 0 0 -
MindSpeed-RL 20 0.679 105.678M 69.116 64 6.636
Async (ours) 20 0.737 123.999M 21.453 48 33.449
The evaluation methodology is identical to that used in the first set of experiments. According to the results, for the 32B-scale model, the ratio of the TPSPD metric between Mindspeed-RL and our asynchronous framework is 1 : 5.05.
@@
Figure 6: TPSPD between MindSpeed-RL and Async (ours)
@@ 6.3 Asynchronous Scalability Experiments
Table 3: Scalability comparison results. TPSPD performance of training Qwen3-8B on 16, 32, and 64 NPUs using our asynchronous framework.
Model NPUs Training Instance Inference Instance Training Steps Tokens/Sec/Device
Qwen3-8B 16 1 4 20 188.162
32 2 8 20 171.824
64 4 16 20 163.208
@@
Figure 7: TPS of our framework under different scales of data parallelism
@@ 7 Conclusion
This paper presents an asynchronous on-policy reinforcement learning algorithm that maintains the same training stability as synchronous RL methods. The feasibility of performing reinforcement learning training with micro-batches is theoretically demonstrated. Prompt-level asynchronous parallelism maximizes the overlap between inference and training. During training, a three-mode (policy, old policy, reference) unified and identically distributed architecture is introduced, along with a group-mask attention mechanism that provides clear advantages when prompts are long and responses are short. Experiments show that the proposed framework achieves a three- to five-fold improvement in end-to-end performance over current mainstream RL training frameworks on the NPU platform. The results also support the equivalence in training effectiveness between the asynchronous and synchronous methods. Furthermore, the decoupling of training and inference allows both to be scaled independently and flexibly, achieving optimal balance and maximizing the speed-to-device ratio. The decoupling of training architecture and algorithm also implies that the proposed framework can be applied to other on-policy reinforcement learning algorithms without any modifications for compatibility with asynchronous training architectures.

### v3

@@ 5 Implementation
Our implementation is built on the PyTorch training framework, using a 3D-parallel distributed architecture. We incorporate parts of Megatron-Core v0.8.0 and DeepSpeed v0.17.2 Rasley et al. (2020). On the NPU platform, we additionally use MindSpeed v2.1.0_core_r0.8.0, torch_npu, and optimized operators such as npu_fusion_attention, an accelerated attention kernel supporting custom masks and serving as a counterpart to flash_attention Dao (2023).
For inference, we use vLLM, with ascend_vLLM 0.9.2.rc1 on NPUs. As in fine-tuning, all RL samples are computed at their native sequence lengths for both forward and backward passes without padding, following dynamic-length training.
@@ 6 Experiments
Framework Training Steps AIME24 Trained Tokens (M) Training Hours NPUs Tokens/Sec/Device
Baseline 0 0.725 0 0 0
MindSpeed-RL 100 0.733 706.355 198.942 16 61.641
Sync (ours) 100 0.746 827.306 143.678 16 99.966
Async (ours) 100 0.758 917.609 82.861 16 192.259
Table 1: Accuracy and training efficiency comparison on Qwen3-8B. All frameworks were trained for 100 steps, and training time and token counts were extracted from logs. For our synchronous and asynchronous methods under the training–inference decoupled architecture, 8 NPUs were used for one training instance ($\text{TP}=8,\text{PP}=1$) and 8 NPUs for four inference instances ($\text{TP}=2$). MindSpeed-RL used all 16 NPUs for two training instances ($\text{TP}=8,\text{PP}=1$) and eight inference instances ($\text{TP}=2$).
Framework Training Steps AIME24 Trained Tokens (M) Training Hours NPUs Tokens/Sec/Device
Baseline 0 0.726 0 0 0
MindSpeed-RL 20 0.696 107.874 70.647 64 6.627
Async (ours) 20 0.737 123.999 21.453 48 33.449
Table 2: Accuracy and training efficiency comparison on DeepSeek-R1-Distill-Qwen-32B. All frameworks were trained for 20 steps, and training time and token counts were taken from logs. Our asynchronous setup uses 32 NPUs for one training instance ($\text{TP}=8,\text{PP}=4$) and 16 NPUs for four inference instances ($\text{TP}=4$). MindSpeed-RL uses all 64 NPUs for two training instances ($\text{TP}=8,\text{PP}=4$) and 16 inference instances ($\text{TP}=4$).
This study conducts comparative experiments on mathematical reasoning tasks, evaluating three reinforcement learning frameworks: MindSpeed-RL, which uses a shared-resource architecture for training and inference; the synchronous reinforcement learning framework implemented in this work under a decoupled training–inference design; and the proposed asynchronous reinforcement learning framework. The main goal is to experimentally compare periodic asynchronous and synchronous training, assess the performance of the proposed system against mainstream synchronous frameworks, and evaluate the scalability of the periodic asynchronous system. The primary evaluation metric is end-to-end training throughput, measured by tokens trained per second per device (TPSPD), calculated as the total number of tokens processed divided by the total elapsed time and the number of devices used. Accuracy metrics are provided for reference to verify algorithm correctness, as no modifications were made to the reinforcement learning algorithm.
Experiments were conducted on air-cooled Ascend-910B NPUs, with each node equipped with eight 64 GB NPUs. The intra-node interconnect bandwidth is 196 GB/s, and the inter-node bandwidth is 100 Gb/s.
@@ 6.1 Model and Environment Configuration
The models compared in this section are Qwen3-8B  Team (2025) and DeepSeek-R1-Distill-Qwen-32B  DeepSeek-AI (2025) (since the current version of Mindspeed-RL does not yet support reinforcement learning training for Qwen3-32B, we instead use DeepSeek-R1-Distill-Qwen-32B as the training model.). The training data comes from the publicly available math problem dataset DeepScaler Luo and others (2025). All reinforcement learning experiments use the GRPO algorithm. The accuracy test set is the AIME24 Zhang and Math-AI (2024).
@@ 6.2 Comparison between Asynchronous and Synchronous Training
Given that the prompts for the samples are relatively short while the responses are longer, Shared-Prompt Attention is disabled in the experiments. In the first set of experiments, Qwen3-8B was used as the training model, with a context length of 16k. As shown in Table 1, in terms of TPSPD, the training speed ratio among the three implementations was $\mathbf{1\!:\!1.62\!:\!3.12}$. The model used in the second set of experiments is DeepSeek-R1-Distill-Qwen-32B, with a context length of 16k, and only two experiments were conducted. As shown in Table 2, the evaluation methodology is identical to that used in the first set of experiments. According to the results, for the 32B-scale model, the ratio of the TPSPD metric between Mindspeed-RL and our asynchronous framework is $\mathbf{1\!:\!5.05}$.
@@
For accuracy evaluation, we employed a rule-based method where the predicted answer from the policy model is considered correct if it can be accurately extracted and matches the ground-truth answer in the dataset; otherwise, it is deemed incorrect. The tests were conducted using the inference en gine vLLM 0.9.1.rc with an output length limit of 32k. Each test sample was evaluated 8 times, with inference parameters set to temperature = 0.7, top_p = 0.95, and top_k = 20. The results show that the models trained by all three frameworks achieved similar accuracy on the AIME24 benchmark, with any differences falling within the expected range of random fluctuations. This supports the equivalence in performance between the asynchronous and synchronous methods implemented in this study.
We also monitored the reward scores during training for all three frameworks, as shown in Figure 5(a). The reward trajectories of our synchronous and asynchronous methods are almost identical, suggesting that the asynchronous approach is computationally equivalent to the synchronous one. Both curves exhibit a slight upward trend, likely because the baseline model was pre-trained on similar data. Figure 5(b) presents a comparison of reward scores between Mindspeed-RL and our asynchronous architecture, and it appears that the trends are generally consistent.
(a) Synchronous vs. Asynchronous.
(b) Mindspeed-RL vs. our Asynchronous.
@@
Figure 5: Training step-wise average reward score comparison.
@@
Figure 6: TPSPD: MindSpeed-RL vs. Async (ours).
@@ 6.3 Asynchronous Scalability Experiments
Model NPUs Training Steps Tokens/Sec/Device
Qwen3-8B 16 20 188.162
32 20 171.824
64 20 163.208
Table 3: Scalability comparison results. TPSPD performance of training Qwen3-8B on 16, 32, and 64 NPUs using our asynchronous framework. The training–to–inference instance ratio is fixed at 1:4.
Figure 7: TPS of our framework under different data-parallel scales.
@@ 7 Conclusion
This paper introduces an asynchronous on-policy reinforcement learning algorithm that maintains the same training stability as synchronous RL methods. The feasibility of performing reinforcement learning training with micro-batches is theoretically demonstrated. Prompt-level asynchronous parallelism is employed to maximize the overlap between inference and training. A unified three-mode (policy, old policy, reference) architecture with identically distributed components is introduced, along with a shared-prompt attention mechanism that offers distinct advantages for long prompts and short responses. Experimental results demonstrate that the proposed framework achieves a three- to five-fold improvement in end-to-end performance over current mainstream RL training frameworks on the NPU platform. These results further support the equivalence in training effectiveness between the asynchronous and synchronous methods. Additionally, the decoupling of training and inference enables independent and flexible scaling, optimizing the balance between speed and device utilization. The decoupling of the training architecture and algorithm also ensures the proposed framework’s compatibility with other on-policy reinforcement learning algorithms, allowing them to benefit from the advantages of asynchronous training without requiring algorithmic modifications, nor ensuring the stability of modified algorithms.